What is the momentum of a 12.2 kg turtle swimming upstream with a speed of 3.3 m/s.



1
Expert's answer
2021-05-27T10:19:14-0400

The magnitude of the momentum vector (of a single moving body) can be found as the total weight of the object times the velocity, thus the momentum of the swimming turtle (expressed in SI units) is:


"p_{turtle}=mv=(12.2\,kg)(3.3\,\frac{m}{s})=40.26\,\frac{kg\cdot m}{s}"


In conclusion, the turtle's momentum is 40.26 kg·m/s.


Reference:

  • Young, H. D., Freedman, R. A., & Ford, A. L. (2006). Sears and Zemansky's university physics (Vol. 1). Pearson education.